a number is multiplied by 5, and then 3 is subtracted from the product. the result is 12. find the number
Answer by luke94(89)

Hi there, firstly consider the number to be variable 'x' and then write an equation of what you are wanting to solve with...
equation: 5x - 3 = 12
now it is just a matter of algebra solving
5x - 3 = 12
5x = 12 + 3
5x = 15
x = 15/5
therefore x = 3
you can then put this number into x if you wish to check that it works out
5(3) - 3 = 15 - 3 = 12
problem solved :)